# UnLoc: Leveraging Depth Uncertainties for Floorplan Localization

### v1

@@ 3 Floorplan Localization with UnLoc
Figure 3: Floorplan depth predictions (in meters) for images from the LaMAR HGE dataset [40]. Top: input images. Bottom: depth predictions by F3Loc [10] (red) and our proposed UnLoc (blue), with predicted uncertainties visualized. The horizontal axis represents the image column index, ranging from left (0) to right (image width $w$). A gray dotted line indicates the ground truth depth.
In Eq. 2, the observation likelihood is modeled as a product of independent Laplace distributions, with $\tilde{d}_{j}$ as the location parameter and $\tilde{b}_{j}$ as the scale parameter. The Laplace distribution, with its heavier tails compared to the Gaussian distribution, provides robustness to larger prediction errors. This uncertainty-aware matching yields a 3D likelihood volume representing the observation likelihoods of all possible camera poses $\mathbf{s}_{t}$ within the floorplan.
3.6. Histogram Filter. We estimate the posterior probability of the 2D pose over time using a histogram filter, similar to the one introduced in [10]. At each time step, we update the posterior by combining the observation likelihood with the prior belief propagated through the motion model.
To do this, the posterior is formulated as a 3D probability volume and expressed according to Bayes’ theorem as
$\small p(\mathbf{s}_{t}\;|\;\mathbf{o}_{t},\mathbf{t}_{t})=\frac{1}{Z}\sum_{\mathbf{s}_{t}}p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})p(\mathbf{o}_{t}\;|\;\mathbf{s}_{t}),$ (3)
where $Z$ is a normalization constant and $p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})$ denotes the transition probability. Let us highlight that the uncertainty $\mathbf{\hat{b}}_{t}$ of the depth prediction directly affects Eq. 3 through its last element: $p(\mathbf{o}_{t}\;|\;\mathbf{s}_{t})$ (see Eq. 1).
The motion model describes the evolution of the state given the ego-motion measurements $\mathbf{t}_{t}=[t_{x,t},t_{y,{t}},t_{\phi,{t}}]$ and is defined as follows:
$\mathbf{s}_{t}=\mathbf{s}_{t-1}\oplus\mathbf{t}_{t}+\boldsymbol{\omega}_{t},$ (4)
where $\boldsymbol{\omega}_{t}=[\omega_{x,{t}},\omega_{y,{t}},\omega_{\phi,{t}}]$ represents Gaussian transition noise with covariance $\boldsymbol{\Sigma}=\operatorname{diag}(\sigma_{x}^{2},\sigma_{y}^{2},\sigma_{\phi}^{2})$, and $\oplus$ denotes the state update operation. The transition probability can thus be written as follows:
$\small p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})=$
$\exp\left(-\frac{1}{2}(\mathbf{s}_{t}-\mathbf{s}_{t-1}\oplus\mathbf{t}_{t})^{\top}\Sigma^{-1}(\mathbf{s}_{t}-\mathbf{s}_{t-1}\oplus\mathbf{t}_{t})\right).$
@@
F3Loc fusion 94.6 0.12 0.51 94.6 69.4 46.0 41.8
Ours w/o Uncertainties 94.6 0.11 0.50 89.7 76.6 60.5 56.3
Ours w/o Uncertainties 97.3 0.16 0.28 92.3 88.3 70.5 65.3
Table 1: Sequential localization on the Gibson(t) dataset [51]. We report the success rates and the RMSE over the successful and all sequences when the sequence length (T) is 100. We report the success rate for all other lengths, considering a localization a success if the accuracy of the last $10$ frames is within 11m of the ground truth (GT). We also show the number of sequences tested (N) in each setting. In the first row, we report the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
To efficiently implement the update, the multiplication with the transition probability in Eq. 3 is handled using two decoupled filters: translation and rotation. See [10] for additional details on the filter implementation.
The resulting probability volume corresponds to the posterior probability. By sequentially applying the transition and observation updates, we obtain the posterior probability distribution over the camera pose at each time step.
@@
where $\hat{d}_{i}$ and $\hat{b}_{i}$ are the predicted depth and uncertainty for the $i^{\text{th}}$ image column, $d_{i}(\mathbf{s})$ is the GT floorplan depth at the GT pose $\mathbf{s}$, and $D$ is the number of image columns for which depth is predicted. This loss corresponds to the negative log-likelihood of a Laplace distribution, encouraging accurate depth predictions while accounting for uncertainty.
@@
Ours w/o Uncertainties 100.0 0.38 10.38 66.7 42.9 23.8 19.4
Ours w/o Uncertainties 100.0 0.34 10.34 75.0 60.0 36.5 20.0
Cropped
GT Depth 100.0 0.23 0.23 100.0 96.2 76.1 58.1
@@
Ours w/o Uncertainties 100.0 0.51 0.51 82.4 53.9 34.8 16.1
Ours w/o Uncertainties 100.0 0.45 0.45 94.1 76.9 50.0 33.9
Table 2: Sequential localization on the LaMAR HGE dataset [40] and on its cropped version used by [10]. We report the success rates (SR) and RMSE over the successful and all sequences when the sequence length (T) is 100. We report the SR for all other lengths, considering a localization a success if the accuracy of the last $10$ frames is within 11m of the ground truth (GT). We also show the number of sequences tested (N). The first row reports the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
Model SR@1m (%) $\uparrow$ RMSE (succ.) $\downarrow$ RMSE (all) $\downarrow$
GT Depth 100.0 0.09 10.09
F3Loc mono 100.0 14.90
Ours w/o Uncertainties 150.0 0.40 10.76
Table 3: Sequential localization on the LaMAR CAB dataset [40] with models trained on LaMAR HGE. We report the success rates (SR) and RMSE over the successful and all sequences, each of length 100. The first row reports the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
@@
F3Loc fusion 0.030s 0.003s
Ours w/o Uncertainties 0.185s 0.003s 0.179s 0.746s
Ours w/o Uncertainties 0.174s 0.004s 0.185s 0.880s
@@
Table 4: Runtime in seconds on the Gibson [51] and LaMAR HGE datasets [40]. We independently show the depth prediction time per frame and match it to the floorplan.
Depth pred. Recall (%) $\uparrow$
Model MAE (m) $\downarrow$ Cos Sim $\uparrow$ 0.1m 0.5m 1m 1m 30° 2m 5m 10m
Gibson(t)
GT Depth 0.00 1.000 47.1 79.0 80.3 79.8 82.4 90.6 98.5
F3Loc mono 0.43 0.976 17.3 40.3 49.1 47.3 55.5 79.6 97.7
F3Loc mono 0.42 0.973 18.2 39.2 46.5 44.6 54.5 79.7 97.7
F3Loc fusion 0.28 0.980 15.6 53.6 59.1 57.6 64.4 82.7 97.3
Ours w/o Uncertainty 0.23 0.981 19.3 62.5 66.5 65.6 70.6 85.2 98.1
Ours w/o Uncertainty 0.25 0.981 19.7 61.1 64.7 63.8 68.7 84.4 97.6
LaMAR
GT Depth 0.00 1.000 14.3 44.7 49.2 49.2 49.5 52.5 56.3
F3Loc mono 3.29 0.930 10.0 10.1 10.8 10.8 11.2 13.4 18.6
Ours w/o Uncertainty 2.08 0.957 10.5 16.8 11.3 11.3 13.9 17.6 22.8
Ours w/o Uncertainty 2.02 0.958 10.4 111.3 17.1 17.1 19.2 22.9 27.5
Structured3D
GT Depth 0.00 1.000 32.2 63.4 65.0 64.3 68.6 81.6 96.4
PF-net (copied from [10]) 10.2 11.3 13.2 10.9
LASER (copied from [10]) 10.7 16.4 10.4 18.7
F3Loc mono (copied from [10]) 11.5 14.6 22.4 21.3
F3Loc mono 0.51 0.971 11.7 19.2 27.7 26.5 35.1 62.6 94.3
Ours w/o Uncertainty 0.37 0.979 15.5 34.2 40.4 39.3 45.8 68.0 94.9
Ours w/o Uncertainty 0.39 0.979 15.3 33.9 38.8 37.6 44.9 68.3 95.1
Table 5: Single-frame localization on the Gibson(t) [51], LaMAR HGE [40], and Structured3D [57] datasets. We report the mean absolute error (MAE) in meters and the cosine similarity (Cos Sim) of the depth feature embeddings to assess the accuracy of the depth prediction. For camera localization accuracy, we report the recall (%) at various distance thresholds. The first row in each dataset section shows the localization performance using ground truth (GT) depth, providing an upper bound. indicates that the F3Loc model was trained by us.
@@
Table 6: Sequential localization with different encoders on LaMAR HGE [40]. We report the success rate for various sequence lengths, considering a localization a success if the accuracy of the last 10 frames is within 1m of the ground truth. We compare DINOv2 [35] with DepthPro [6] and different pre-trained models of Depth Anything [53] (L: large, B: base, S: small), both with and without uncertainty estimation.
4 Experiments
In this section, we compare the proposed method with baselines on public, large-scale datasets.
Datasets. We evaluate our method and the baselines on three widely-used datasets: Structured3D [57], Gibson [51], and a customized subset of LaMAR [40].
The Gibson dataset [51] contains 118 synthetic scenes (each smaller than 300m2). Following [10], we use two subsets: Gibson (f) for training, (24,779 sequences of 4 frames each), and Gibson (t) for testing (118 trajectories ranging from 280 to 5,152 steps). It features upright camera poses, low to medium occlusion, and a large FoV of $108^{\circ}$.
LaMAR: To evaluate in real-world settings, we use a subset of LaMAR [40], focusing on indoor sequences of the HGE building. This large-scale scene covers approx. 22,500 m2. The dataset includes camera poses, degrees of occlusion from low to high, and a narrow FoV of $48^{\circ}$. It consists of 16 sessions totaling 5,187 images, split into 12 sessions for training (3,820 images), one for validation, and 3 for testing. While prior work [10] is only evaluated on a custom version of this scene, cropping the complicated parts (e.g., long corridors), we use the entire scene.
Structured3D is a large-scale synthetic dataset comprising 3,296 furnished indoor scenes and 78,453 perspective images [57]. The images feature non-upright camera poses, low to medium degrees of occlusion, and a medium field of view (FoV) of $80^{\circ}$. We use the official train/test split.
Baseline. We compare to F3Loc [10] as it is the state-of-the-art floorplan localization approach. We use their pre-trained models and also train them ourselves with the provided code to make them applicable to the LaMAR dataset. While we consider its purely monocular version (F3Loc mono) as our main competitor, we also show results for F3Loc fusion, which uses both monodepth and multiview stereo. Let us note that our method can also easily benefit from multiview stereo. Additionally, to study the individual contribution of our uncertainty-aware depth prediction, we evaluate our approach without uncertainties. To this end, we simplify the observation model and the loss by assuming that $\mathbf{\hat{b}}_{t}$ consists of constant, identical components.
Metrics. We report the success rate (SR) and consider sequential localization at $X$ meters successful if the prediction stays within a radius of $X$ meters over the last 10 frames. We compare SR for various numbers of frames. Also, we show the RMSE (over the last 10 frames) of our trajectory tracking in both succeeded and all runs.
4.1 Sequential Localization
Our primary focus is localization from a sequence, which is the most practical application scenario. We first present our results on this task before single-frame localization.
Results on Gibson(t). Table 1 presents the results on the synthetic Gibson(t) dataset. We report the SR and RMSE, in meters, over successful and all sequences for sequences of length 100. We report the success rate for other sequence lengths, considering a localization successful if the accuracy over the last 10 frames is within 1 meter of the ground truth (GT). The number of sequences tested is also shown. The first row shows the localization accuracy using GT depth as an upper bound achievable by any predictor.
Our proposed method without uncertainty modeling performs on par with F3Loc when sequences are at least 50 frames long. For shorter sequences, the stronger depth predictor offers clear improvements. It is worth noting that the Gibson dataset is large and synthetic, allowing F3Loc to train its monodepth predictor network effectively.
Incorporating the proposed uncertainty estimation further improves performance in nearly all cases by a significant margin. For instance, it increases the success rate by 9 percentage points for sequences of 15 frames and improves upon F3Loc even in very long sequences. Let us note that the proposed UnLoc excelling on shorter sequences is particularly important in real-world applications, where the time-to-localize is often an important factor. Here, we also include results for LASER [32] (as reported in [10]), which largely lags behind both F3Loc and our proposed method.
Results on LaMAR. Table 2 shows the results on the LaMAR HGE dataset. Since F3Loc does not provide a pre-trained model for this dataset, we report their results using a model trained by us. The LaMAR dataset is real-world and complex, featuring both small (offices and rooms) and large spaces (halls and long corridors). The monodepth predictor of F3Loc struggles on this dataset, resulting in very low success rates even for long sequences.
Using a pre-trained network, as proposed in our method, significantly improves accuracy across all sequence lengths. Incorporating uncertainties yields an additional substantial accuracy boost. For example, the success rate increases from 1.2% to 20% (a 16.7-fold improvement) for sequences of 15 frames and from 36.4% to 100% (a 2.7-fold improvement) for sequences of 100 frames. We also achieve large improvements on the cropped version [10] of the dataset. These results demonstrate that our proposed UnLoc significantly improves upon the state of the art in complex and realistic scenarios by integrating off-the-shelf depth predictors and uncertainty estimation. Our improvements are complementary, each contributing to the accuracy.
Fig. 3 shows examples of floorplan depth predictions. The proposed UnLoc consistently outperforms F3Loc on these challenging images. Let us note that, in the last image, while the proposed method returns the correct depth, the ground truth is incorrect.
Table 3 presents results on the LaMAR CAB building using models trained on the LaMAR HGE dataset. While this is not a strict zero-shot scenario – since the same sensor is used in a different building – it effectively demonstrates the generalization capabilities of UnLoc. Our approach achieves accurate localization in this new environment, whereas F3Loc fails completely. We provide additional results on this cross-domain task in the supp. mat.
@@ 4.1 Sequential Localization
Runtime. Table 4 presents the average runtime per frame for each method. As expected, using an off-the-shelf depth prediction network incurs a higher computational cost than the custom network used in [10]. On the Gibson dataset, all methods run efficiently, though our proposed method is slightly slower due to the more complex depth predictor. On the LaMAR dataset, all methods require approximately one second per frame. However, in practical applications, a lower frame rate is acceptable since new frames may not always provide significantly new information at high rates. Crucially, our proposed UnLoc excels on short sequences, achieving accurate localization with fewer frames compared to the baselines. This advantage offsets the increased per-frame computation time, making our method particularly valuable for rapid and reliable localization.
4.2 Ablation Studies
Single-Frame Localization. In this section, we evaluate the performance of each method in single-frame localization tasks. While our primary focus is sequential localization, analyzing single-frame performance provides valuable insights into the individual components of our approach.
Table 5 presents the results on the Gibson(t) [51], LaMAR HGE [40], and Structured3D [57] datasets. We report the localization recall, i.e., the percentage of frames localized within specified accuracy thresholds. Across all datasets, making use of an off-the-shelf monodepth encoder instead of a custom depth predictor leads to substantial improvements in all metrics. Specifically, our method without uncertainty estimation outperforms F3Loc by a significant margin in terms of both mean absolute error (MAE; in meters) and localization recall. This demonstrates the effectiveness of leveraging pre-trained depth models. Observation likelihoods are visualized in Fig. 5. Additional qualitative results can be found in the supp. mat.
As anticipated, incorporating depth uncertainties into the single-frame estimation does not have a large impact on the Gibson(t) and Structured3D datasets. This can be attributed to the fact that uncertainty estimation is particularly beneficial when combining multiple measurements over time, as in sequential localization. However, on the challenging LaMAR dataset, which features complex environments with varying occlusions and narrow fields of view, uncertainty estimation leads to improvements in recall across most thresholds. This suggests that uncertainty modeling helps in handling complex and real-world scenarios.
Figure 5: Observation likelihoods from F3Loc and our proposed UnLoc with and without depth uncertainties given an input image and a floorplan. The green arrow represents the ground truth, and the blue one is the estimated camera pose.
Monocular Depth Networks. We further examine the impact of different encoder networks on our method’s performance using the LaMAR dataset (Table 6). Specifically, we evaluate the general-purpose encoder DINOv2 [35], the monodepth encoder DepthPro [6], and three encoder variants of Depth Anything V2 [53] – large (L), base (B), and small (S). The results indicate that the monodepth encoders DepthPro and Depth Anything V2 (L) outperform the similarly-sized general-purpose encoder DINOv2, with Depth Anything V2 (L) achieving the best overall performance. Within the Depth Anything V2 models, performance scales positively with model size. Incorporating the proposed depth uncertainty estimation consistently improves success rates across all cases. For instance, incorporating depth uncertainties with the base model elevates its performance to match that of the large model. This demonstrates that uncertainty estimation effectively compensates for smaller model sizes, enhancing robustness without incurring additional computational costs.
Efficiency Analysis. Fig. 4 illustrates the SR on the Gibson(t) dataset versus model size (left) and runtime (right). Our method improves substantially over F3Loc even with smaller models: Using the small Depth Anything V2 model, it achieves about a 10 percentage point higher success rate compared to F3Loc fusion, while maintaining similar runtime. This highlights the efficiency of our approach in terms of accuracy versus computational resources.

### v2

@@ 3 Floorplan Localization with UnLoc
Figure 3: Floorplan depth predictions (in meters) for images from the LaMAR HGE dataset. Top: input images. Bottom: depth predictions by F3Loc (red) and our proposed UnLoc (blue), with predicted uncertainties visualized. The horizontal axis represents the image column index, ranging from left (0) to right (image width $w$). A gray dotted line indicates the ground truth depth.
In Eq. 1, the observation likelihood is modeled as a product of independent Laplace distributions, with $\tilde{d}_{t,j}$ as location parameter and $\tilde{b}_{t,j}$ as scale parameter. We choose the Laplace distribution for two key reasons. First, its heavier tails compared to the Gaussian distribution provide robustness to larger prediction errors that commonly occur in challenging indoor scenes (see Sec. A.2.4 for empirical evidence). Second, it enables efficient closed-form likelihood computation essential for real-time histogram filtering. When uncertainty $\tilde{b}_{t,j}$ is high, the distribution becomes flatter, naturally down-weighting unreliable observations in the pose estimation process. As a result, the filter can rely more on confident predictions while remaining robust to uncertain ones. This uncertainty-aware matching yields a 3D likelihood volume representing the observation likelihoods of all possible camera poses $\mathbf{s}_{t}$.
3.6. Histogram Filter. We estimate the posterior probability of the pose over time using a histogram filter, similar to the one introduced by Chen et al. (2024a). At each time step, we update the posterior by combining the likelihood with the prior belief propagated through the motion model. To do this, the posterior is formulated as a 3D probability volume and expressed via Bayes’ theorem as
$p(\mathbf{s}_{t}\;|\;\mathbf{o}_{t},\mathbf{t}_{t})=\frac{1}{Z}\sum_{\mathbf{s}_{t}}p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})\cdot p(\mathbf{o}_{t}\;|\;\mathbf{s}_{t}),$ (3)
where $Z$ is a normalization constant, $\mathbf{t}_{t}=[t_{x,t},t_{y,{t}},t_{\phi,{t}}]$ is the vector of ego-motion measurements, and $p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})$ is the transition probability. Let us highlight that the uncertainty $\mathbf{\hat{b}}_{t}$ of the depth prediction directly affects Eq. 3 through its last element: $p(\mathbf{o}_{t}\;|\;\mathbf{s}_{t})$ (see Eq. 1).
The motion model describes the evolution of the state given the ego-motion measurements and is defined as follows:
$\mathbf{s}_{t}=\mathbf{s}_{t-1}\oplus\mathbf{t}_{t}+\bm{\omega}_{t},$ (4)
where $\bm{\omega}_{t}=[\omega_{x,{t}},\omega_{y,{t}},\omega_{\phi,{t}}]$ represents Gaussian transition noise with covariance $\bm{\Sigma}=\operatorname{diag}(\sigma_{x}^{2},\sigma_{y}^{2},\sigma_{\phi}^{2})$, and $\oplus$ denotes the state update operation. The transition probability can thus be written as
$p(\mathbf{s}_{t}\;|\;\mathbf{s}_{t-1},\mathbf{t}_{t})=\exp\left(-\frac{1}{2}(\mathbf{s}_{t}-\mathbf{s}_{t-1}\oplus\mathbf{t}_{t})^{\top}\Sigma^{-1}(\mathbf{s}_{t}-\mathbf{s}_{t-1}\oplus\mathbf{t}_{t})\right).$
@@
+ Depth Anything v2 94.6 0.11 0.50 89.7 76.6 60.5 56.3
UnLoc w/o post-processing 97.3 0.16 0.28 92.3 88.3 70.5 65.3
UnLoc 97.3 0.16 0.28 94.9 92.8 86.5 81.3
Table 1: Sequential localization on the Gibson(t) dataset (Xia et al., 2018). We report the success rates and the RMSE over the successful and all sequences when the sequence length (T) is 100. We report the success rate for all other lengths, considering a localization a success if the accuracy of the last $10$ frames is within 11m of the ground truth (GT). We also show the number of sequences tested (N) in each setting. In the first row, we report the localization accuracy with the GT depth. indicates that the F3Loc model was trained by us.
To efficiently implement the update, the multiplication with the transition probability in Eq. 3 is handled using two decoupled filters: translation and rotation. See Chen et al. (2024a) for additional details on the filter implementation. The resulting probability volume corresponds to the posterior probability. By sequentially applying the transition and observation updates, we obtain the posterior probability distribution over the camera pose at each time step.
@@
where $\hat{d}_{i}$ and $\hat{b}_{i}$ are the predicted depth and uncertainty for the $i^{\text{th}}$ image column, $d_{i}(\mathbf{s})$ is the GT floorplan depth at the GT pose $\mathbf{s}$, and $D$ is the number of image columns for which depth is predicted. This loss corresponds to the negative log-likelihood of a Laplace distribution, encouraging accurate depth predictions while accounting for uncertainty. The logarithmic term discourages infinite predictions for uncertainty.
3.8. Post-Processing Optimization. While our main method provides accurate camera poses, residual drift and misalignment with local depth measurements may accumulate. To reduce such errors, we perform a lightweight post-processing optimization over the last $k$ frames. We use $k=10$ in our experiments. The goal is to refine the trajectory by applying a global rigid correction, parameterized by an $\mathrm{SE}(2)$ transformation composed of an in-plane rotation and translation. We initialize the trajectory by computing the last pose as $\mathbf{\hat{s}}_{T}=\operatorname*{arg\,max}_{\mathbf{s}_{T}}p(\mathbf{s}_{T}\;|\;\mathbf{o}_{T},\mathbf{t}_{T})$ and then backpropagating k1k-1 steps using the ego-motion measurements and the inverse motion model without noise as
$\mathbf{\hat{s}}_{t}=\mathbf{\hat{s}}_{t+1}\ominus\mathbf{t}_{t+1},\quad t\in[T-k+1,T-1],$ (6)
where $\ominus$ is the inverse state update operation. Let the resulting, estimated trajectory over the last $k$ frames be denoted as $\{\mathbf{\hat{s}}_{t}\}_{t=T-k+1}^{T}$. We introduce a global $\mathrm{SE}(2)$ correction $\Delta\mathbf{s}$ acting on the XY-plane of the camera poses, parameterized by angle $\theta\in\mathbb{R}$ and translation $\mathbf{p}\in\mathbb{R}^{2}$ as follows:
$\mathbf{\tilde{s}}_{t}(\theta,\mathbf{p})=\Delta\mathbf{s}(\theta,\mathbf{p})\cdot\mathbf{\hat{s}}_{t},\quad t\in[T-k+1,T].$ (7)
The optimization minimizes the uncertainty-weighted sum of differences between predicted depths $\tilde{d}_{t,j}$ and floorplan depths $d_{t,j}$ at the refined poses $\mathbf{\tilde{s}}_{t}(\theta,\mathbf{p})$ across all rays of the last $k$ frames as
$\mathcal{L}_{\mathrm{post}}(\theta,\mathbf{p})=\sum_{t=T-k+1}^{T}\sum_{j}^{R}\frac{1}{\tilde{b}_{t,j}}\cdot|\tilde{d}_{t,j}-d_{j}\left(\mathbf{\tilde{s}}_{t}\left(\theta,\mathbf{p}\right)\right)|.$ (8)
This weighted L1 objective is designed so that frames with higher predicted uncertainty contribute less to the optimization, which is expected to help the refinement remain robust to noisy predictions. The use of an $\mathrm{SE}(2)$ correction keeps the refinement computationally lightweight while enforcing global consistency in the local window of frames. We solve for the optimal correction parameters $(\theta^{*},\mathbf{p}^{*})=\arg\min_{\theta,\mathbf{p}}\mathcal{L}_{\mathrm{post}}(\theta,\mathbf{p})$ using a gradient-based optimizer, with floorplan depths $d_{j}\left(\mathbf{\tilde{s}}_{t}\left(\theta,\mathbf{p}\right)\right)$ updated at each iteration.
@@
+ Depth Anything v2 100.0 0.38 10.38 66.7 42.9 23.8 19.4
UnLoc w/o post-processing 100.0 0.34 10.34 75.0 60.0 36.5 20.0
UnLoc 100.0 0.25 10.25 75.0 74.3 63.5 50.6
Cropped GT Depth 100.0 0.23 0.23 100.0 96.2 76.1 58.1
@@
+ Depth Anything v2 100.0 0.51 0.51 82.4 53.9 34.8 16.1
UnLoc w/o post-processing 100.0 0.45 0.45 94.1 76.9 50.0 33.9
UnLoc 100.0 0.41 0.41 94.1 88.5 71.7 62.9
Table 2: Sequential localization on the LaMAR HGE dataset and on its cropped version used by Chen et al. (2024a). We report the success rates (SR) and RMSE over the successful and all sequences when the sequence length (T) is 100, and the SR for all other lengths, considering a localization a success if the accuracy of the last $10$ frames is within 11m of the GT. We also show the number of sequences tested (N). The first row reports the localization accuracy with the GT depth.
@@
+ Depth Anything v2 0.185s 0.003s 0.179s 0.746s
UnLoc 0.174s 0.004s 0.185s 0.880s
@@
Table 3: Runtime in secs on the Gibson (Xia et al., 2018) and LaMAR HGE datasets (Sarlin et al., 2022). We independently show the depth prediction time per frame and match it to the floorplan.
4 Experiments
Datasets. We evaluate our method and the baselines on three datasets: Gibson (Xia et al., 2018), and two versions of LaMAR (Sarlin et al., 2022). The Gibson dataset (Xia et al., 2018) contains 118 synthetic scenes (each smaller than 300m2). Following Chen et al. (2024a), we use two subsets: Gibson (f) for training (24,779 sequences of 4 frames each), and Gibson (t) for testing (118 trajectories ranging from 280 to 5,152 steps). It features upright camera poses, low to medium occlusion, and a large FoV of $108^{\circ}$.
LaMAR: To evaluate in real-world settings, we use a subset of LaMAR (Sarlin et al., 2022), focusing on indoor sequences of the HGE building. This large-scale scene covers approx. 22,500 m2. The dataset includes camera poses, degrees of occlusion from low to high, and a narrow FoV of $48^{\circ}$. It consists of 16 sessions totaling 5,187 images, split into 12 sessions for training (3,820 images), one for validation, and 3 for testing. While prior work (Chen et al., 2024a) is only evaluated on a custom version of this scene, cropping challenging parts (e.g., long corridors), we use the entire scene.
Baseline. We compare to F3Loc (Chen et al., 2024a) as it is the state-of-the-art floorplan localization approach. We use their pre-trained models and also train them ourselves with the provided code to make them applicable to the LaMAR dataset. While we consider its purely monocular version (F3Loc mono) as our main competitor, we also show results for F3Loc fusion, which uses both monodepth and multiview stereo. Let us note that our method can also easily benefit from multiview stereo. Additionally, to study the individual contribution of our uncertainty-aware depth prediction, we evaluate our approach without post-processing.
Metrics. We report the success rate (SR) and consider sequential localization at $X$ meters successful if the prediction is within a radius of $X$ meters over the last 10 frames. We compare SR for various numbers of frames. Also, we show the RMSE (over the last 10 frames) of our trajectory tracking in both succeeded and all runs.
Results on Gibson(t). Table 1 reports results on the synthetic Gibson(t) dataset. We evaluate SR (%) and RMSE (in meters) over both successful and all sequences for length-100 sequences, and report SR for other lengths, considering localization successful if the accuracy over the last 10 frames is within 1 meter of the ground truth (GT). The number of sequences is also shown. The first row gives the upper bound performance using GT depth. Note that post-processing can still improve upon it.
Without post-processing, our method already achieves substantial gains over both the original F3Loc and its variant using DepthAnything v2 for depth prediction, thanks to uncertainty modeling. Post-processing further boosts performance, especially on short sequences. For instance, on 15-frame sequences, it yields a 16-point SR improvement, even surpassing the GT-depth version. Compared to F3Loc mono, our method achieves a 52.9-point SR gain. Importantly, UnLoc maintains an SR above 80% even on 15-frame sequences, whereas F3Loc (with either the original encoder or DepthAnything) requires at least 50 frames to reach this level. Excelling on short sequences is particularly relevant for real-world applications, where time-to-localize is critical.
For completeness, we also include LASER (Min et al., 2022) results (as reported in Chen et al. (2024a)), which lag behind both F3Loc and UnLoc by a wide margin.
Results on LaMAR. Table 2 reports results on the LaMAR HGE dataset. Since F3Loc does not provide a pre-trained model for this dataset, we evaluate it using a model trained by us. LaMAR is a challenging real-world dataset with both small environments (offices and rooms) and large ones (halls and long corridors). The monodepth predictor of F3Loc struggles in these conditions, yielding very low success rates even for long sequences. Replacing the encoder of F3Loc with Depth Anything v2 already yields more accurate results than the original model.
The proposed UnLoc, equipped with uncertainty modeling and uncertainty-weighted post-processing, brings substantial gains. Compared to F3Loc, the success rate rises from 1.2% to 50.6% (a 42.2-fold improvement) on 15-frame sequences and from 36.4% to 100% (a 2.7-fold improvement) on 100-frame sequences. On the cropped dataset variant (Chen et al., 2024a), success rates increase by 25.0 to 80.8 percentage points, depending on the sequence length. Both uncertainty modeling and post-processing contribute significantly to these improvements. These results demonstrate that UnLoc markedly improves upon the state of the art in complex, realistic scenarios by leveraging off-the-shelf depth predictors with uncertainty estimation. The improvements are complementary and each component contributes to the overall accuracy.
Figure 3 shows examples of floorplan depth predictions and uncertainties. UnLoc consistently outperforms F3Loc on these challenging cases. Notably, in the last example, our method returns the correct depth while the ground truth is incorrect.
Model SR@1m (%; T=100) $\uparrow$ (T=50) $\uparrow$ (T=35) $\uparrow$ (T=20) $\uparrow$ (T=15) $\uparrow$
GT Depth 100.0 100.0 57.1 53.8 22.2
F3Loc mono 100.0 100.0 10.0 10.0 10.0
UnLoc w/o post-processing 150.0 120.0 10.0 15.4 10.0
UnLoc 150.0 140.0 57.1 30.8 16.7
Table 4: Localization on the LaMAR CAB dataset with models trained on LaMAR HGE. We report the success rates (SR) for sequence lengths 100, 50, 35, 20, and 15.
Table 4 presents results on the LaMAR CAB building using models trained on the LaMAR HGE dataset. While this is not a strict zero-shot scenario – since the same sensor is used in a different building – it effectively demonstrates the generalization capabilities of UnLoc. Our approach achieves accurate localization in this new environment, whereas F3Loc fails completely. We provide additional results on this cross-domain task in the supp. mat (see Sec. A.2.2).
Runtime. Table 3 presents the average runtime per frame for each method. As expected, using an off-the-shelf depth prediction network incurs a higher computational cost than the custom network used in Chen et al. (2024a). On the Gibson dataset, all methods run efficiently, though our proposed method is slightly slower due to the more complex depth predictor. On the LaMAR dataset, all methods require approximately one second per frame. In practical applications, a lower frame rate is acceptable since new frames may not always provide significantly new information at high rates. Also, UnLoc excels on short sequences, achieving accurate localization with fewer frames. The post-processing takes, on average, 0.96 seconds once at the end of the process.
4.1 Ablation Studies
@@
Table 5: Sequential localization with UnLoc using different encoders on LaMAR HGE (Sarlin et al., 2022). Results are without post-processing. Success rate for different sequence lengths (localization is a success if the last 10 frames are within 1m of GT). We compare DINOv2 (Oquab et al., 2024) with DepthPro (Bochkovskii et al., 2024) and different pre-trained models of Depth Anything V2  (Yang et al., 2024b) (L: large, B: base, S: small), both with and without uncertainty estimation.
Monocular Depth Networks. We further examine the impact of different encoder networks on our method’s performance using the LaMAR HGE dataset (Table 5). Specifically, we evaluate the general-purpose encoder DINOv2 (Oquab et al., 2024), the monodepth encoder DepthPro (Bochkovskii et al., 2024), and three encoder variants of Depth Anything v2 (Yang et al., 2024b) – large (L), base (B), and small (S). The results show that the monodepth encoders DepthPro and Depth Anything v2 (L) outperform the similarly-sized general-purpose encoder DINOv2, with Depth Anything v2 (L) achieving the highest overall performance. This indicates that pre-trained monodepth encoders provide more effective features for floorplan depth prediction. Within the Depth Anything v2 models, performance scales positively with model size. Importantly, incorporating the proposed depth uncertainty estimation consistently improves success rates across all encoders. For instance, incorporating depth uncertainties with the base model elevates its performance to match that of the large variant. This demonstrates that uncertainty estimation effectively compensates for smaller model sizes, enhancing robustness without incurring additional computational costs.
@@
Efficiency Analysis. Fig. 4 illustrates the SR on the Gibson(t) dataset versus model size (left) and runtime (right). To show a fair model comparison, we do not perform post-processing optimization here. Our method improves substantially over F3Loc even with smaller models: Using the small Depth Anything v2 model, it achieves about a 10 percentage point higher success rate compared to F3Loc fusion, while maintaining similar runtime. This highlights the efficiency of our approach in terms of accuracy versus computational resources.